Suppose that (X1, X2,,,,Xn) are iid random variables. Find the maximum likelihood estimator of theta for the following distributions

1) Poi(theta)

2) N(Mu, theta)

3) Exp(theta)
Answer #1

(1)

The pdf of Poisson distribution is

P(X =x) =-피

Here we have

P(X=x_{1})=rac{e^{- heta} heta^{x_{1}}}{x_{1}!}

PlX =x2) = _.ra!

.

.

.

P(X=x_{n})=rac{e^{- heta} heta^{x_{n}}}{x_{n}!}

So the likelihood function will be

L( heta)=P(X_{1}=x_{1})P(X_{2}=x_{2})...P(X_{n}=x_{n})=e^{-n heta}rac{lambda^{sum_{i=1}^{n}x_{i}}}{prod_{i=1}^{n}x_{i}!}

Taking log of both sides gives

lnleft [L( heta) ight ]=lnleft [e^{-n heta}rac{ heta^{sum_{i=1}^{n}x_{i}}}{prod_{i=1}^{n}x_{i}!} ight ]=-n heta+ sum_{i=1}^{n} x_{i} ln heta- ln left ( prod_{i=1}^{n}x_{i}! ight )

Differentiating both sides gives

rac{d}{d heta}lnleft [L( heta) ight ]=-n+rac{sum_{i=1}^{n}x_{i}}{ heta}

  

Equating this equal to zero gives

-n+rac{sum_{i=1}^{n}x_{i}}{ heta}=0

heta=rac{1}{n} sum_{i=1}^{n}x_{i}

That is required MLE is

ilde{ heta}=rac{1}{n} sum_{i=1}^{n}x_{i}

(2)

The pdf of normal distribution is

f(x) =-1

Let X1, X2, ...Xn is a random sample from the normal distribution. So we have

f(x_{1})=rac{1}{sqrt{2pi heta^{2}}}e^{-rac{(x_{1}-mu)^{2}}{2 heta^{2}}}

f(x_{2})=rac{1}{sqrt{2pi heta^{2}}}e^{-rac{(x_{2}-mu)^{2}}{2 heta^{2}}}

.

.

.

f(x_{n})=rac{1}{sqrt{2pi heta^{2}}}e^{-rac{(x_{n}-mu)^{2}}{2 heta^{2}}}

The likelihood function is

L(mu, heta^{2}:x_{1},x_{2},...,x_{n})=f(x_{1})f(x_{2})...f(x_{n})=left [rac{1}{sqrt{2pi heta^{2}}} ight ]^{n}cdot e^{-rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{2 heta^{2}}}

Taken log of both sides gives

l(mu, heta^{2}:x_{1},x_{2},...,x_{n})=-rac{n}{2}ln(2pi)-rac{n}{2}ln( heta^{2})- rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{2 heta^{2}}

Differentiating above with respect to mu gives

rac{partial }{partial mu}l(mu, heta^{2}:x_{1},x_{2},...,x_{n})=rac{sum_{i=1}^{n}(x_{i}-mu)}{ heta^{2}}

Equating it to zero gives

rac{sum_{i=1}^{n}(x_{i}-mu)}{ heta^{2}}=0

sum_{i=1}^{n}(x_{i}-mu)=0

sum_{i=1}^{n}x_{i}-nmu=0

hat{mu}=rac{1}{n}sum_{i=1}^{n}x_{i}=ar{x}

-------------------------

Differentiating above with respect to sigma gives

rac{partial }{partial heta^{2}}l(mu,sigma^{2}:x_{1},x_{2},...,x_{n})=-rac{n}{2 heta^{2}}+rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{2 heta^{4}}

Equating it to zero gives

-rac{n}{2 heta^{2}}+rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{2 heta^{4}}=0

rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{2 heta^{4}}=rac{n}{2 heta^{2}}

rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{ heta^{2}}=n

heta^{2}=rac{sum_{i=1}^{n}(x_{i}-mu)^{2}}{n}

heta^{2}=rac{1}{n}left [ sum_{i=1}^{n}(x^{2}_{i}+mu^{2}-2mu x_{i}) ight ]=rac{1}{n}left [ sum_{i=1}^{n}x^{2}_{i}+nmu^{2}-2mu sum_{i=1}^{n}x_{i} ight ]=left [rac{1}{n} sum_{i=1}^{n}x^{2}_{i}+mu^{2}-2mu ar{x} ight ]

Putting estimator of my gives

heta^{2}=left [rac{1}{n} sum_{i=1}^{n}x^{2}_{i}+ar{x}^{2}-2ar{x}^{2} ight ]=rac{1}{n} sum_{i=1}^{n}x^{2}_{i}-ar{x}^{2}

So required estimate is

hat{ heta}^{2}=left [rac{1}{n} sum_{i=1}^{n}x^{2}_{i}+ar{x}^{2}-2ar{x}^{2} ight ]=rac{1}{n} sum_{i=1}^{n}x^{2}_{i}-ar{x}^{2}

(3)

From the definition of pdf we have

f(x_{1})=rac{1}{ heta} e^{-x_{1}/ heta}

f(r2)

.

.

.

f(x_{n})=rac{1}{ heta} e^{-x_{n}/ heta}

So the likelihood function will be

-T2 エ1 Taking log of both sides gives

lnleft [L( heta) ight ]=lnleft [rac{1}{ heta^{n}} e^{-(x_{1}+x_{2}+...x_{n})/ heta} ight ]=ln(1)-nln( heta)-rac{sum_{i=1}^{n}x_{i}}{ heta}

Differentiating above with respect to heta gives

rac{d}{d heta}lnleft [L( heta) ight ]=-rac{n}{ heta}+rac{sum_{i=1}^{n}x_{i}}{ heta^{2}}

Equating above to zero gives

-rac{n}{ heta}+rac{sum_{i=1}^{n}x_{i}}{ heta^{2}}=0

heta=rac{sum_{i=1}^{n} x_{i}}{n}=ar{x}

Hence, required estimate is

ilde{ heta}=rac{sum_{i=1}^{n} x_{i}}{n}=ar{x}.